Grade 11Mechanics

A thin sheet of metal of uniform thickness is cut into the shape bounded by the linex=a and y=kx^2. Find the coordinates of the centre of mass.

riya
9 Years agoGrade 11

1 Answer

Rituraj Tiwari
5 Years ago

To find the coordinates of the center of mass of a thin sheet of metal cut into the shape bounded by the lines \(x = a\) and the curve \(y = kx^2\), we need to apply the concept of centroid for a two-dimensional area. The center of mass (or centroid) of a shape can be computed by finding the average position of all points in the shape, weighted by area. The formula for the coordinates of the centroid \((\bar{x}, \bar{y})\) of a region \(R\) is given by:\[\bar{x} = \frac{1}{A} \int_R x \, dA \]\[\bar{y} = \frac{1}{A} \int_R y \, dA \]where \(A\) is the area of the region \(R\).Let's break this down step by step.

Identifying the Shape and Boundaries

The shape is defined by the parabola \(y = kx^2\) and the vertical line \(x = a\). The area we are interested in lies between these two boundaries. The other boundary is the x-axis, where \(y = 0\). Thus, our region of integration is bounded by \(x = 0\) to \(x = a\) and \(y = 0\) to \(y = kx^2\).

Calculating the Area (A)

The area \(A\) can be calculated using the integral of the function from \(x = 0\) to \(x = a\):\[A = \int_0^a kx^2 \, dx\]Calculating this integral, we have:\[A = k \left[ \frac{x^3}{3} \right]_0^a = k \frac{a^3}{3}\]This gives us the total area of the shape.

Finding the Coordinates of the Centroid

Now, let's find \(\bar{x}\) and \(\bar{y}\).

Calculating \(\bar{x}\)

Using the formula for \(\bar{x}\):\[\bar{x} = \frac{1}{A} \int_0^a x \cdot kx^2 \, dx = \frac{1}{A} \int_0^a kx^3 \, dx\]Calculating the integral:\[\int_0^a kx^3 \, dx = k \left[ \frac{x^4}{4} \right]_0^a = k \frac{a^4}{4}\]Thus,\[\bar{x} = \frac{1}{A} \cdot k \frac{a^4}{4} = \frac{4}{ka^3} \cdot k \frac{a^4}{4} = \frac{a}{4}\]

Calculating \(\bar{y}\)

Next, we find \(\bar{y}\):\[\bar{y} = \frac{1}{A} \int_0^a \frac{1}{2} y^2 \, dx\]To find this, we need to calculate:\[\bar{y} = \frac{1}{A} \int_0^a \frac{1}{2} (kx^2)^2 \, dx = \frac{1}{A} \int_0^a \frac{k^2 x^4}{2} \, dx\]Calculating the integral:\[\int_0^a \frac{k^2 x^4}{2} \, dx = \frac{k^2}{2} \left[ \frac{x^5}{5} \right]_0^a = \frac{k^2 a^5}{10}\]Thus,\[\bar{y} = \frac{1}{A} \cdot \frac{k^2 a^5}{10} = \frac{1}{\frac{ka^3}{3}} \cdot \frac{k^2 a^5}{10} = \frac{3k a^5}{10ka^3} = \frac{3a^2}{10}\]

Final Coordinates of the Center of Mass

Putting this all together, the coordinates of the center of mass of the thin sheet of metal are:\[\left(\bar{x}, \bar{y}\right) = \left(\frac{a}{4}, \frac{3a^2}{10}\right)\]This process demonstrates how calculus can be used to find the center of mass of a complex shape by breaking it down into manageable steps. If you have any further questions about this topic or need clarification on any part, feel free to ask!